OpenStudy (anonymous):

for some reason i keep getting the wrong answer for this fraction: simplify. (11/20 divided by 55/3) - (11/20 times 55/3) this is the only one i keep missing for some reason.

OpenStudy (studygurl14):

Remember that when you divide a fraction, you multiply by its reciprocal \(\Large (\frac{11}{20}\div \frac{55}{3})-(\frac{11}{20}\times\frac{55}{3})=(\frac{11}{20}\times\frac{3}{55})-(\frac{11}{20}\times\frac{55}{3})=?\)

OpenStudy (anonymous):

yes i have that part down

OpenStudy (studygurl14):

Okay, so what do you get?

OpenStudy (anonymous):

so far i got 33/1100 - 605/60

zepdrix (zepdrix):

Things work out better when you don't multiply stuff together :)

OpenStudy (studygurl14):

Factor out the 11/20

OpenStudy (studygurl14):

Is that where you're headed @zepdrix ?

zepdrix (zepdrix):

From your last step, I probably would have done a quick cancellation before moving forward.\[\large\rm \frac{\color{orangered}{11}\cdot3}{20\cdot\color{orangered}{55}}-\frac{11\cdot55}{20\cdot3}\]These orange pieces have something in common, yes?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Yes, factoring works nicely as well :) I don't think she's up to that point though.

zepdrix (zepdrix):

Cancellation gives us,\[\large\rm \frac{\cancel{\color{orangered}{11}}\cdot3}{20\cdot\cancel{\color{orangered}{55}}5}-\frac{11\cdot55}{20\cdot3}\]

zepdrix (zepdrix):

\[\large\rm \frac{3}{20\cdot5}-\frac{11\cdot55}{20\cdot3}\]I would STILL recommend not multiplying things out. It's easier to see what each denominator is missing from this setup. The first fraction is missing a 3, while the second fraction is missing a 5, ya?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{\frac33}\cdot\frac{3}{20\cdot5}-\frac{11\cdot55}{20\cdot3}\cdot\color{royalblue}{\frac55}\]Which gives us something like,\[\large\rm \frac{3\cdot3}{20\cdot3\cdot5}-\frac{11\cdot55\cdot5}{20\cdot3\cdot5}\]Which we can combine into a single fraction, right?\[\large\rm \frac{3\cdot3-11\cdot55\cdot5}{20\cdot3\cdot5}\]And then it's probably safe to multiply things out from this point :) And simplify.

OpenStudy (anonymous):

When we do bigger fractions that involve more operations is it easier to do it the way we just did it ?

zepdrix (zepdrix):

Maybe not always. If you look at my last post, that is a big mess of numbers, right? If it becomes too much to carry around like that, then I suppose you might reach a point where simplifying things down BEFORE proceeding forward makes more sense. Though, leaving the denominator expanded out as I did will almost always better more beneficial. It just makes it way easier to see what each denominator is "missing".

OpenStudy (anonymous):

makes sense. i know a lot of times when i try to work the whole thing out long ways i get confused by the numbers because it turns to a long process sometimes

zepdrix (zepdrix):

hehe ya :)) try to get comfortable with AT LEAST leaving the bottoms expanded out. math is tricky! but keep at it! work hard! study! be blessed! :D

OpenStudy (anonymous):

aww thank you i been trying my best and so far its doing me good !

zepdrix (zepdrix):

good good good c:

OpenStudy (anonymous):

thank you for helping me tonight :)

zepdrix (zepdrix):

np